Soluve using the multiplication principle first. Then use the elimination method:
7p+5q=2
8p-9q=17
I know a common number needs to be found but how do I choose which one?
Thanks!

Answer by Alan3354(69443)

7p+5q=2
8p-9q=17
---------
You can eliminate either p or q. In this case, it's 7 & 8 or 5 & 9.
I'll do 5 & 9
7p+5q=2 --> 63p + 45q = 18
8p-9q=17--> 40p - 45q = 85
-----------
63p + 45q = 18
40p - 45q = 85
-------------- Add
103p = 103
p = 1
------
7p+5q=2
7 + 5q = 2
5q = -5
q = -1